OpenStudy (anonymous):

Calc III Help Lagrange Multipliers Maximize: f(x,y)= sqrt(6-x^2-y^2) Constraint: x+y-2=0

OpenStudy (turingtest):

first step is to get the system of equations got that yet?

OpenStudy (anonymous):

The first partials?

OpenStudy (turingtest):

get the gradient of f and set the components equal to a constant multiple of the gradient of the constraint then solve for the constant, lambda

OpenStudy (anonymous):

I got the gradient f = -1,-1 is that correct? Also when I was solving for lambda I am getting -1 and it that doesnt seem right.

OpenStudy (turingtest):

how did you get a number for the gradient? what is the partial of f wrt x ?

OpenStudy (anonymous):

Fx = -x / sqrt(6-x^2-y^2)

OpenStudy (anonymous):

Isn't the grad F equal to the coeficients in what you are trying to min/max?

OpenStudy (turingtest):

no the grad is just \[\nabla f(x,y)=<f_x,f_y>\]left as variables

OpenStudy (anonymous):

Alright. So after finding grad f. I am to find the partial for the constraint correct?

OpenStudy (anonymous):

partials**

OpenStudy (turingtest):

right, you need\[\nabla f=\lambda\nabla g\]where g is the constrain then set each component equal\[f_x=\lambda g_x\]\[f_y=\lambda g_y\]then solve the system

OpenStudy (anonymous):

both Gx and Gy are 1 correct?

OpenStudy (turingtest):

right

OpenStudy (anonymous):

So I am then just setting the Fx=Fy and solving?

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

Alright and once I am able to get either everything to equal x or y I would pulg that back into either Fx or Fy which ever is easier to solve?

OpenStudy (anonymous):

Or into the constrain?

OpenStudy (anonymous):

constraint*

OpenStudy (turingtest):

no, you plug it into the constraint what result did you get from your system?

OpenStudy (anonymous):

\[y = (x \sqrt{6-x ^{2}-y ^{2}})\div \sqrt{6-x ^{2}-y ^{2}}\]

OpenStudy (turingtest):

I fear you have over-complicated this...

OpenStudy (anonymous):

Damn

OpenStudy (turingtest):

\[f_x=fy\]\[\frac{-x}{\sqrt{6-x^2-y^2}}=\frac{-y}{\sqrt{6-x^2-y^2}}\]\[x=y\]

OpenStudy (anonymous):

Alright. So after putting that back into the constraint x=1 and y=1 correct?

OpenStudy (turingtest):

right

OpenStudy (anonymous):

Thank you very much for clearing that up. So from the start find Fx, Fy, Gx, and Gy then set Fx= Lambda Gx and Fy= Lambda Gy then solve for lambda in 1 of the 2 equations plug it into the other solve for either x or y then plug that number into the constraint then solve for x and y in the contraint and that will be your max or min.

OpenStudy (turingtest):

yes there could, of course, be three variables at times also note that lambda always drops out of the problem, so you never actually need to find its value also, don't forget to check that whatever point you have actually is a min or max i.e. check the actual value(s) of f you get from the plug-in(s) you're welcome, and good luck!

OpenStudy (turingtest):

my hobby-horse reference for much of this kind of thing is this page http://tutorial.math.lamar.edu/Classes/CalcIII/LagrangeMultipliers.aspx

OpenStudy (anonymous):

Where I check the max/min I would plug those into the f(x) correct?

OpenStudy (turingtest):

exactly

OpenStudy (anonymous):

Gotchya. Alright thank you again for the help.

OpenStudy (turingtest):

anytime!

OpenStudy (anonymous):

Quick question when doing extremas does D=FxxFyy-(Fxy)^2

OpenStudy (turingtest):

yes it does :D